# Zhixuan Chu1∗♠, Yan Wang1∗, Longfei Li1, Zhibo Wang2, Zhan Qin2, Kui Ren2

### v1

@@ 3.2 The Real Situation
This bias originates from the biases in the pre-training data. For example, consider the semantic prompt “I … talking about weddings”. Due to biases present in the pre-training data, the LLM may have learned to associate weddings with positive sentiments such as love, happiness, and celebration. As a result, the semantic direction representation $R^{cd}$ learned from this prompt may implicitly suggest a positive direction (e.g., "love") for the output, even if the explicit steering prompt is not provided. As shown in Figure 1, this edge can be visually observed experimentally. The implicit influence of the semantic direction representation $R^{cd}$ on the output direction can be problematic when attempting to steer the LLM’s output towards a desired attribute. If the steering representation $\Delta R$ is not independent of $R^{cd}$, the resulting output may be biased by the inherent associations learned during pre-training, leading to suboptimal or unintended results. To ensure unbiased steering of the output, it is crucial to disentangle the steering representation $\Delta R$ from the semantic direction representation $R^{cd}$.
3.3 Causal Analysis of Our Solutions
Based on the causal analysis presented in the previous sections, we propose a solution called LLMGuardaril to address the bias introduced by the semantic direction representation $R^{cd}$ in the steering process. As shown in Figure 3 (c), in addition to constructing pairs of prompts (positive and negative steering prompts with the same semantic prompt), we need to block the edge from the semantic direction representation $R^{cd}$ of the semantic prompt to the direction representation $R_{+}/R_{-}$. By doing so, the direction representation is only influenced by the steering prompt $S_{+}/S_{-}$, enabling us to obtain an unbiased steering representation $\Delta R$ through steering engineering. This approach aligns with the desired hypothetical situation.
We utilize adversarial learning to remove this confounding bias [58; 59]. The objective is to achieve debiasing of the influence $R^{cd}$ on the direction representation $R_{+}/R_{-}$ during the adversarial learning process. The adversarial learning process includes two main components, i.e., prediction reconstruction learning and debias learning. In prediction reconstruction learning, the goal is to ensure that the original semantic information remains unchanged during the debiasing process. This is achieved by minimizing the prediction reconstruction loss, which measures the cross-entropy between the original output and the output generated.
@@ 3.3 Causal Analysis of Our Solutions
Figure 4: The framework of LLMGuardaril, which is a plug-and-play algorithmic framework designed to obtain the unbiased steering representation for LLMs while seamlessly integrating with their existing architecture.
4 Methodology
4.1 Intervened Layer Selection
Previous studies have investigated the information encoded in different layers of transformer-based models. Tenney et al. [60] found that earlier layers in BERT encode lower-level information, such as part-of-speech tags, while later layers capture more semantic information. Similarly, Zou et al. [18] and Li et al. [20] have observed that the optimal intervention layers for steering the model’s output vary depending on the specific task and desired attribute.
@@ 4.1 Intervened Layer Selection
Empirically, we find that the middle layers of the pre-trained language model tend to be the most effective for intervention. This observation aligns with previous findings [18; 20] suggesting that the middle layers capture a balance between low-level syntactic information and high-level semantic information, making them suitable for steering the model’s output towards the desired attribute or concept. By selecting the intervened layers based on probing accuracy, we can focus the intervention on the most relevant layers, thereby improving the efficiency and effectiveness of the steering process. The selected layers $L^{*}$ are then used in the Debias LoRA Block to obtain the debiased representations and calculate the steering representation.
4.2 Unbiased Steering Representations
@@
4.2.1 Debias Training Framework
As shown in Figure 4, LLMGuardaril is a plug-and-play algorithmic framework designed to obtain unbiased steering representations for LLMs while seamlessly integrating with their existing architecture. The framework consists of two key components: the Debias LoRA Block and the Domain Probing module. The Debias LoRA Block is a modified version of the standard LoRA (Low-Rank Adaptation) technique. Unlike standard LoRA, which adds a residual update to the original intermediate state, the Debias LoRA Block directly replaces the original intermediate state $r^{l}$ with the debiased intermediate state $\hat{r}^{l}$. This is achieved through the following operation:
$\hat{r}^{l}=\Delta Wr^{l-1}=BAr^{l-1}$ (1)
@@ 4.2.1 Debias Training Framework
4.2.2 Adversarial Learning
The training process of LLMGuardaril involves adversarial learning, where the Debias LoRA Block and the Domain Probing module are optimized simultaneously. The objective is to use the Domain Probing module to debias the influence of $R^{cd}$ on the direction representations $R_{+}$ and $R_{-}$ during the adversarial learning process [61].
@@ 4.2.2 Adversarial Learning
The adversarial learning process consists of two main components: prediction reconstruction learning and debias learning. In the debias learning, the objective is to debias the representation through the Domain Probing module. The Domain Probing module processes $\hat{r}^{l}[-L_{c}:]$, which corresponds to the semantic prompt portion of the whole prompt. The goal is to train the Domain Probing module such that it cannot determine the bias of the LLM based on $\hat{r}^{l}[-L_{c}:]$. To facilitate this, a Gradient Reversal Layer (GRL) [62] is introduced in the debias learning process. The debias loss is defined as follows:
$\mathcal{L}_{debias}=\sum^{N}_{i=1}\Bigl{(}y^{direction}-\text{GradRev}(f(\hat{r}^{l}[-L_{c}:],\eta))\Bigl{)}$ (2)
@@
In prediction reconstruction learning, the objective is to ensure that the original semantic information remains unchanged during the debiasing process. This is achieved by minimizing the prediction reconstruction loss, which measures the cross-entropy between the original output $y^{output}$ by original framework $\phi$ using original representation $R$ and the output generated by LLMGuardaril $\hat{\phi}$ using the debiased representations $\hat{R}$:
$\mathcal{L}_{pre}=\text{CEloss}(y^{output},\hat{\phi}(\hat{R}))$ (3)
@@
$\mathcal{L}=\mathcal{L}_{pre}+\alpha\mathcal{L}_{debias}$ (4)
@@
By employing this adversarial learning framework, LLMGuardaril enables the extraction of steering representations that are less influenced by the inherent biases in the LLM’s pre-training data. The debiased intermediate representations generated by the Debias LoRA Block can then be used to guide the LLM’s output toward the desired attributes or concepts while mitigating the impact of unwanted biases.
4.2.3 Obtaining the Steering Representation
@@
    $\Delta\hat{r}^{*}=\frac{1}{N}\sum_{j=1}^{N}\frac{1}{n}\sum_{i=1}^{n}\Delta\hat{r}_{ij}^{*},$ (5)
@@
The resulting $\Delta\hat{r}^{*}$ represents the final steering representation obtained from the debias training process. This steering representation encodes the desired steering direction, taking into account the contrast between positive and negative steering prompts, averaged across tokens and samples. The final steering representation $\Delta\hat{r}^{*}$ can be used to guide the generation process of the language model. By adding this steering representation to the intermediate representations of the LLM during inference, we can steer the model’s output towards the desired attributes or concepts while mitigating the influence of biases present in the pre-training data.
4.3 Explanation of Output
@@
$\text{Similarity}_{i}=\overline{\Delta\hat{r}^{*}}\cdot\overline{r^{*}_{i}},$ (6)
@@
$\text{Alignment}=\frac{1}{n}\sum_{i=1}^{n}\text{Similarity}_{i}$ (7)
@@
4.4 Control of Output
The steering representations are injected into the model and activated during inference, directing the model’s response toward the desired direction. Unlike most existing methods [20; 63; 35; 34] that simply add or subtract a constant steering representation regardless of the token representation, we propose a more sophisticated approach to control the output. Our method takes into account the relationship between the generated token representation and the steering representation, allowing for more fine-grained and context-aware control of the output.
To establish a connection between the generated token representation and the steering representation, we employ a projection operation inspired by [18]. This operation amplifies the component of the token representation that aligns with the steering representation, effectively emphasizing the desired direction in the output. Let $r^{*}_{i}\in\mathbb{R}^{K\times d}$ denote the representation of the generated token $i$ obtained from the intervened layers. This is achieved by projecting out the component in the direction of steering representation $\Delta\hat{r}^{*}$, and the operation can be defined as
$\hat{r}^{*}_{i}=r^{*}_{i}+\beta\times\frac{r^{*\mathsf{T}}_{i}\Delta\hat{r}^{*}}{\|\Delta\hat{r}^{*}\|^{2}}\Delta\hat{r}^{*},$ (8)
@@ 4.4 Control of Output
Furthermore, the projection operation can be extended to incorporate multiple steering representations simultaneously. In scenarios where we want to steer the output towards multiple desired directions, we can compute the projections onto each steering representation separately and combine them using appropriate weighting schemes. This enables a more comprehensive control over the output, allowing us to incorporate multiple desired attributes or concepts.
@@
5 Experiments
We evaluate the performance of our proposed LLMGuardaril framework on four key attributes that serve as guardrails for large language models: truthfulness, toxicity, bias, and harmfulness. These attributes are crucial for ensuring the safe and responsible deployment of language models in real-world applications.
5.1 Baselines
In this section, we introduce a diverse set of baseline methods that aim to steer the behavior of large language models towards desired attributes or concepts. By comparing LLMGuardaril against these baselines, we can assess its effectiveness in achieving the desired steering while maintaining the model’s general knowledge and capabilities.
Base model (Few-shot) [64] is an in-context learning method that does not require fine-tuning of the language model. This technique leverages a small set of high-quality samples to construct prompts, which are then used to guide the Language Learning Model (LLM) in generating answers that closely resemble the provided examples. To implement this approach, we first create a set of 3-5 queries. These queries are then fed into GPT-4, a state-of-the-art language model, to generate high-quality answers. The resulting query-answer pairs serve as in-distribution examples, which are subsequently used to construct prompts during the inference process. By providing the LLM with these carefully crafted prompts, the model is able to generate answers that are more aligned with the desired output, thereby improving the overall quality of the answers.
Linear Artificial Tomography (LAT-Reading) [18] is a representation reading technique that aims to locate emergent representations for high-level concepts and functions within a network. The LAT pipeline consists of three key steps: designing stimulus and task, collecting neural activity, and constructing a linear model. A stimulus set and task template are carefully designed to elicit distinct neural activity for the target concept or function. Next, neural activity is collected from specific token positions in the model’s hidden states. Finally, a linear model, such as PCA or a supervised method, is applied to the collected neural activity to identify directions that align with the target concept or function. The resulting “reading vectors” can be generally used to steer the output direction.
LAT-Contrast [18] is a representation control technique that builds upon the reading vectors obtained through LAT. During inference, the same input is run through the model using a pair of contrastive prompts, producing two different representations. The difference between these representations forms a Contrast Vector, which serves as a stimulus-dependent controller. The Contrast Vectors are used to modify the model’s representations at each layer using operations such as linear combination, piece-wise operations, or projection. Compared to the above LAT-Reading, it requires over 3×3$\times$3 × more inference compute for the positive prompt, negative prompt, and the original prompt.
Low-Rank Representation Adaptation (LoRRA) [18] is another representation control technique that addresses the computational overhead associated with calculating Contrast Vectors during inference. LoRRA involves fine-tuning low-rank adapter matrices connected to the model’s attention weights using a specific loss function applied to representations. The loss function is designed to guide the model’s representations towards the desired target representations, which can be obtained using methods such as LAT-Contrast. During training, the adapters are updated to minimize the difference between the model’s current representations and the target representations. Once trained, the adapters are merged into the model, resulting in modified representations without additional computational burden during inference. LoRRA provides an efficient way to control the model by directly optimizing the model’s representations.
Activation Addition (ActAdd) [34] is a lightweight approach for controlling the behavior of pre-trained language models without fine-tuning or optimization. ActAdd operates by modifying the activations of a frozen model at inference time to steer the output in a desired direction, which is specified through natural language prompts. The method computes the difference in activations resulting from a pair of prompts that represent the desired attribute and its opposite, then adds this delta to the activations at a specific layer during inference. This steers the model’s output towards exhibiting the desired attribute while preserving its general knowledge and capabilities on unrelated tasks.
Mean-Centring [35] is a method for steering the behavior of pre-trained language models by modifying their activations at inference time. The key idea is to extract a “distillation vector” that captures the properties of a target dataset exhibiting the desired behavior. This is done by averaging the activations of the model when processing the target dataset and then subtracting the mean activations of the model on a general training dataset. This mean-centering step removes the model’s general activation bias and isolates the specific direction corresponding to the target behavior.
Contrast-Consistent Search (CCS) [26] is an unsupervised method for discovering latent knowledge in the internal activations of pre-trained language models. Given a set of yes-no questions, CCS constructs “contrast pairs” by answering each question as both “yes” and “no”. It then extracts the model’s hidden representations for these contrast pairs, normalizes them to remove positional biases, and learns a linear mapping from the representations to probabilities of being true or false.
5.2 Benchmarks and Evaluation Metrics
To assess the effectiveness of LLMGuardaril in steering language models towards desired attributes, we conduct experiments on several public datasets that focus on different aspects of content safety and bias [63]. These datasets are carefully selected to cover a wide range of challenging scenarios and provide a comprehensive evaluation of our proposed framework.
[…]
Toxicity. To evaluate the effectiveness of LLMGuardaril in mitigating toxicity, we employ the ToxiGen dataset [66], which contains implicitly toxic and benign sentences mentioning 13 minority groups. We use a revised version of the dataset [67] that reduces noise by filtering out prompts with disagreement among annotators regarding the target demographic group. Through stratified sampling, we select a representative subset of 700 examples for our experiments. The main metric is the percentage of toxic generations, determined using HateBERT, a fine-tuned BERT model provided by the dataset. Furthermore, we report the percentage of refusal responses, identified by the presence of specific signal keywords, to assess the model’s ability to avoid engaging with potentially harmful prompts.
Bias. The BOLD benchmark [68] is employed to evaluate bias in generated responses. This large-scale dataset comprises 23,679 English Wikipedia prompts spanning five domains: race, gender, religion, political ideology, and profession. To manage experiment costs, we sample 120 prompts from each domain. The VADER sentiment score [69] serves as the primary metric, quantifying the sentiment directed towards the population mentioned in each prompt. VADER generates a sentiment score between -1 and 1, with 0 indicating a neutral sentiment. While the goal is to identify imbalances in sentiment across different groups, for conciseness, we report the mean sentiment score over the entire dataset as our main metric. Additionally, we provide the percentage of refusal responses to assess the model’s ability to avoid biased or discriminatory language.
Harmfulness. To evaluate LLMGuardaril’s performance in mitigating harmful content, we utilize the AdvBench dataset [70], which contains 500 harmful behaviors and instructions reflecting toxic, discriminatory, or cybercrime-related actions. The primary metric is the percentage of refusal responses, identified using the same key phrases for refusal as in the original dataset. Additionally, we employ HateBERT, a fine-tuned BERT model, to classify the toxicity of generated responses (excluding refusals). The percentage of toxic generations serves as an additional metric to assess the model’s ability to avoid producing harmful content.
By conducting experiments on these diverse benchmarks, we aim to provide a comprehensive evaluation of LLMGuardaril’s effectiveness in steering language models towards desired attributes while mitigating undesirable behaviors. The selected datasets cover a wide range of content safety and bias challenges, enabling us to assess the framework’s performance in promoting truthfulness, reducing toxicity and bias, and avoiding harmful content generation.
Figure 5: The examples of the prefix steering prompt sets, and the original and intervened outputs by our LLMGuardaril with explainable shading.
5.3 Experiment Settings
Prompt Design. To thoroughly assess the efficacy of LLMGuardaril, we employ a prompt structure that combines a prefix steering prompt, which can be either positive or negative, with an identical semantic prompt. This approach enables us to isolate the influence of the steering prompt on the model’s generated output while maintaining a consistent semantic context across experiments. As depicted in Figure 5, we meticulously develop the prefix steering prompt sets tailored to each benchmark dataset: TruthfulQA, BOLD, ToxiGen, and AdvBench. These prompts are strategically designed to steer the model towards generating content that aligns with the desired attributes or concepts specific to each dataset, such as truthfulness, lack of bias, non-toxicity, and avoidance of harmful content.
Model Selection. We evaluate the guardrail performance of LLMGuardaril using two prominent families of instruction-tuned large language models: Llama2 [71] and Vicuna-V1.5 [72]. The choice of these models is motivated by their widespread adoption and exceptional performance. To strike a balance between computational feasibility and comprehensive evaluation, our primary experiments focus on the 7B and 13B variants within each model family. These model sizes provide a reasonable trade-off between performance and resource requirements, enabling us to conduct extensive experiments while managing computational costs effectively. To gain a deeper understanding of LLMGuardaril’s scalability and its potential to generalize across different model sizes, we expand our experiments to encompass the entire spectrum of model variants within the Vicuna family, including Vicuna-33B.
@@ 5.3 Experiment Settings
Table 1: Performance comparison on four benchmark datasets. $\uparrow$ means higher is better and $\downarrow$ means lower is better.
@@
LLMGuardaril (ours) 48.22 96.77 46.67 88.85 10.15 8.64 0.899 80.96 19.04
5.4 Main Results
Table 1 presents a comprehensive performance comparison of LLMGuardaril against various baseline methods across four benchmark datasets: TruthfulQA, ToxiGen, BOLD, and AdvBench. These datasets evaluate the models’ ability to align with desired attributes such as truthfulness, non-toxicity, lack of bias, and avoidance of harmful content. Across all datasets and model variants (Vicuna-7b, Llama2-7b, and Llama2-13b), LLMGuardaril consistently outperforms the baseline methods. The results highlight LLMGuardaril’s state-of-the-art performance in steering the language models towards the desired attributes. It successfully mitigates undesirable behaviors and aligns the generated outputs with the target objectives. The superior performance of LLMGuardaril compared to the baseline methods highlights the significance of our proposed framework and its potential for real-world applications. As depicted in Figure 5, we provide examples comparing the original LLM output without any guardrails to the intervened output generated by our LLMGuardaril. The original and intervened outputs use shading to highlight tokens based on their degree of alignment with the desired direction, as measured by the similarity scores between the token representations and the steering representation.
Table 2: Ablation studies of our LLMGuardaril model for the key components. $\uparrow$ means higher is better and $\downarrow$ means lower is better.

### v2

@@ 3.2. The Real Situation
This bias originates from the biases in the pre-training data. For example, consider the semantic prompt “I … talking about weddings”. Due to biases present in the pre-training data, the LLM may have learned to associate weddings with positive sentiments such as love, happiness, and celebration. As a result, the semantic direction representation $R^{cd}$ learned from this prompt may implicitly suggest a positive direction (e.g., “love”) for the output, even if the explicit steering prompt is not provided. As shown in Figure 1, this edge can be visually observed experimentally. The implicit influence of the semantic direction representation $R^{cd}$ on the output direction can be problematic when attempting to steer the LLM’s output toward a desired attribute. If the steering representation $\Delta R$ is not independent of $R^{cd}$, the resulting output may be biased by the inherent associations learned during pre-training, leading to suboptimal or unintended results. To ensure unbiased steering of the output, it is crucial to disentangle the steering representation $\Delta R$ from the semantic direction representation $R^{cd}$.
3.3. Causal Analysis of Our Solutions
Based on the causal analysis presented in the previous sections, we propose a solution called LLMGuardrail to address the bias introduced by the semantic direction representation $R^{cd}$ in the steering process. As shown in Figure 2 (c), in addition to constructing pairs of prompts (positive and negative steering prompts with the same semantic prompt), we need to block the edge from the semantic direction representation $R^{cd}$ of the semantic prompt to the direction representation $R_{+}/R_{-}$. By doing so, the direction representation is only influenced by the steering prompt $S_{+}/S_{-}$, enabling us to obtain an unbiased steering representation $\Delta R$ through steering engineering. This approach aligns with the desired hypothetical situation.
We utilize adversarial learning to remove this confounding bias. The objective is to achieve debiasing of the influence $R^{cd}$ on the direction representation $R_{+}/R_{-}$ during the adversarial learning process. The adversarial learning process is to ensure that the original semantic information remains unchanged during the debiasing process. This is achieved by minimizing the prediction reconstruction loss, which measures the cross-entropy between the original output and the output generated.
@@ 3.3. Causal Analysis of Our Solutions
Figure 4. The framework of LLMGuardrail, which is a plug-and-play algorithmic framework designed to obtain the unbiased steering representation for LLMs while seamlessly integrating with their existing architecture. It consists of (1) Intervened Layer Selection: selecting layers based on probing accuracy. (2) Debias Training Framework: including a Debias LoRA Block that replaces the original intermediate state with the debiased intermediate state, and a Domain Probing module implemented as a multi-layer perceptron (MLP) for adversarial learning. The training process optimizes both prediction reconstruction loss $\mathcal{L}_{pre}$ and debias loss $\mathcal{L}_{debias}$. (3) Inference Stage: applying the learned unbiased steering representations to control the LLM’s output using a projection operation.
4. Methodology
4.1. Intervened Layer Selection
Previous studies have investigated the information encoded in different layers of transformer-based models. Tenney et al. (Tenney et al., 2019) found that earlier layers in BERT encode lower-level information, such as part-of-speech tags, while later layers capture more semantic information. Similarly, Zou et al. (Zou et al., 2023a) and Li et al. (Li et al., 2024) have observed that the optimal intervention layers for steering the model’s output vary depending on the specific task and desired attribute.
@@ 4.1. Intervened Layer Selection
Empirically, we find that the middle layers of the pre-trained language model tend to be the most effective for intervention. This observation aligns with previous findings (Zou et al., 2023a; Li et al., 2024) suggesting that the middle layers capture a balance between low-level syntactic information and high-level semantic information, making them suitable for steering the model’s output toward the desired attribute or concept. By selecting the intervened layers based on probing accuracy, we can focus the intervention on the most relevant layers, thereby improving the efficiency and effectiveness of the steering process. The selected layers $L^{*}$ are then used in the Debias LoRA Block to obtain the debiased representations and calculate the steering representation.
4.2. Unbiased Steering Representations
@@
4.2.1. Debias Training Framework
As shown in Figure 4, LLMGuardrail is a plug-and-play algorithmic framework designed to obtain unbiased steering representations for LLMs while seamlessly integrating with their existing architecture. The framework consists of two key components: the Debias LoRA Block and the Domain Probing module. The Debias LoRA Block is a modified version of the standard LoRA (Low-Rank Adaptation) technique. Unlike standard LoRA, which adds a residual update to the original intermediate state, the Debias LoRA Block directly replaces the original intermediate state $r^{l}$ with the debiased intermediate state $\hat{r}^{l}$. This is achieved through the following operation:
(1) $\hat{r}^{l}=\Delta Wr^{l-1}=BAr^{l-1}$
@@ 4.2.1. Debias Training Framework
4.2.2. Debiasing Training
The debiasing training process of LLMGuardrail involves adversarial learning (Chu et al., 2022, 2021, 2024a), where the Debias LoRA Block and the Domain Probing module are optimized simultaneously. The objective is to debias the influence of $R^{cd}$ on the direction representations $R_{+}$ and $R_{-}$ during the adversarial learning process.
@@ 4.2.2. Debiasing Training
The adversarial learning process is to debias the representation through the Domain Probing module. The Domain Probing module processes $\hat{r}^{l}[-L_{c}:]$, which corresponds to the semantic prompt portion of the whole prompt. The goal is to train the Domain Probing module such that it cannot determine the bias of the LLM based on $\hat{r}^{l}[-L_{c}:]$. To facilitate this, a Gradient Reversal Layer (GRL) (Ganin and Lempitsky, 2015) is introduced. The debias loss is defined as follows:
(2) $\mathcal{L}_{debias}=\sum^{N}_{i=1}\Bigl{(}y^{direction}-\text{GradRev}(f(\hat{r}^{l}[-L_{c}:],\eta))\Bigl{)}$
@@
In addition to debiasing, we also need to ensure that the original semantic information remains unchanged during the debiasing process. This is achieved by minimizing the prediction reconstruction loss, which measures the cross-entropy between the original output $y^{output}$ by original framework $\phi$ using original representation $R$ and the output generated by LLMGuardrail $\hat{\phi}$ using the debiased representations $\hat{R}$:
(3) $\mathcal{L}_{pre}=\text{CEloss}(y^{output},\hat{\phi}(\hat{R}))$
@@
(4) $\mathcal{L}=\mathcal{L}_{pre}+\alpha\mathcal{L}_{debias}$
@@
By employing this adversarial learning framework, LLMGuardrail enables the extraction of steering representations that are less influenced by the inherent biases in the LLM’s pre-training data. The debiased intermediate representations generated by the Debias LoRA Block can then be used to guide the LLM’s output toward the desired attributes or concepts while mitigating the impact of unwanted biases.
4.2.3. Obtaining the Steering Representation
@@
    (5) $\Delta\hat{r}^{*}=\frac{1}{N}\sum_{j=1}^{N}\frac{1}{n}\sum_{i=1}^{n}\Delta\hat{r}_{ij}^{*},$
@@
The resulting $\Delta\hat{r}^{*}$ represents the final steering representation obtained from the debias training process. This steering representation encodes the desired steering direction, taking into account the contrast between positive and negative steering prompts, averaged across tokens and samples. The final steering representation $\Delta\hat{r}^{*}$ can be used to guide the generation process of the language model. By adding this steering representation to the intermediate representations of the LLM during inference, we can steer the model’s output toward the desired attributes or concepts while mitigating the influence of biases present in the pre-training data.
4.3. Explanation of Output
@@
(6) $\text{Similarity}_{i}=\overline{\Delta\hat{r}^{*}}\cdot\overline{r^{*}_{i}},$
@@
(7) $\text{Alignment}=\frac{1}{n}\sum_{i=1}^{n}\text{Similarity}_{i}$
@@
4.4. Control of Output
The steering representations are injected into the model and activated during inference, directing the model’s response toward the desired direction. Unlike most existing methods (Li et al., 2024; Wang and Shu, 2023; Jorgensen et al., 2023; Turner et al., 2023) that simply add or subtract a constant steering representation regardless of the token representation, we propose a more sophisticated approach to control the output. Our method takes into account the relationship between the generated token representation and the steering representation, allowing for more fine-grained and context-aware control of the output.
To establish a connection between the generated token representation and the steering representation, we employ a projection operation inspired by (Zou et al., 2023a). This operation amplifies the component of the token representation that aligns with the steering representation, effectively emphasizing the desired direction in the output. Let $r^{*}_{i}\in\mathbb{R}^{K\times d}$ denote the representation of the generated token $i$ obtained from the intervened layers. This is achieved by projecting out the component in the direction of steering representation $\Delta\hat{r}^{*}$, and the operation can be defined as
(8) $\hat{r}^{*}_{i}=r^{*}_{i}+\beta\times\frac{r^{*\mathsf{T}}_{i}\Delta\hat{r}^{*}}{\|\Delta\hat{r}^{*}\|^{2}}\Delta\hat{r}^{*},$
@@ 4.4. Control of Output
Furthermore, the projection operation can be extended to incorporate multiple steering representations simultaneously. In scenarios where we want to steer the output toward multiple desired directions, we can compute the projections onto each steering representation separately and combine them using appropriate weighting schemes. This enables a more comprehensive control over the output, allowing us to incorporate multiple desired attributes or concepts.
@@
Table 1. Performance comparison on four benchmark datasets. $\uparrow$ means higher is better and $\downarrow$ means lower is better.
@@
LLMGuardrail (ours) 48.22 96.77 46.67 88.85 10.15 8.64 0.899 80.96 19.04
5. Experiments
We evaluate the performance of our proposed LLMGuardrail framework on four key attributes that serve as guardrails for large language models: truthfulness, toxicity, bias, and harmfulness. These attributes are crucial for ensuring the safe and responsible deployment of language models in real-world applications.
5.1. Baselines
In this section, we introduce a diverse set of baseline methods that aim to steer the behavior of large language models toward desired attributes or concepts. By comparing LLMGuardrail against these baselines, we can assess its effectiveness in achieving the desired steering while maintaining the model’s general knowledge and capabilities.
Base model (Few-shot) (Brown et al., 2020) is an in-context learning method that does not require fine-tuning of the language model. Linear Artificial Tomography (LAT-Reading) (Zou et al., 2023a) is a representation reading technique that aims to locate emergent representations for high-level concepts and functions within a network. LAT-Contrast (Zou et al., 2023a) is a representation control technique that builds upon the reading vectors obtained through LAT. Low-Rank Representation Adaptation (LoRRA) (Zou et al., 2023a) is another representation control technique that addresses the computational overhead associated with calculating Contrast Vectors during inference. Activation Addition (ActAdd) (Turner et al., 2023) is a lightweight approach for controlling the behavior of pre-trained language models without fine-tuning or optimization. Mean-Centring (Jorgensen et al., 2023) is a method for steering the behavior of pre-trained language models by modifying their activations at inference time. Contrast-Consistent Search (CCS) (Burns et al., 2022) is an unsupervised method for discovering latent knowledge in the internal activations of pre-trained language models.
Table 2. Ablation studies of our LLMGuardrail model for the key components. $\uparrow$ means higher is better and $\downarrow$ means lower is better.